"define a variable, write an equation, and solve the problem"
find 2 consecutive even integers whose sum is 126.
Could you help me solve this problem please I have 6 such problems and don't understand the process with which I'm suppose to solve them. Thanks

Answer by psbhowmick(878)

Even nos. are of the form 2n where n = any non-zero integer.
If the 1st no. is 2n then the second one is obviously 2(n+1).

Their sum is: 2n + 2(n+1) = 126
4n + 2 = 126
4n = 126 - 2
4n = 124
n = 124/4 = 31

So, the nos. are 2x31 and 2(31+1) i.e. 62 and 64 respectively.